write a quadratic equation in the variable x having the given numbers as solutions. type the equation in standard form, ax^2+bx+c=0, where a,b, and c are intergers(no fractions). Solutions -3/8,5/8.
Answer by ewatrrr(24785)


Hi
quadratic equation with solutions -3/8,5/8
(x+3/8)(x -5/8) = 0
x^2 -2/8x - 15/64 = 0 |Multiplying thru by 64 so as alll denominators = 1
64x^2 - 16x - 15 = 0